OpenStudy (anonymous):

Convert the rectangular coordinates to polar coordinates with r > 0 and 0 ≤ θ < 2π. (-squareroot of 6, -squareroot of 2)

OpenStudy (anonymous):

\[\theta = \arctan (-\sqrt{2}/-\sqrt{6}), r = \sqrt{6 + 2}\]

OpenStudy (anonymous):

what is the radian for the \[\theta\]?

OpenStudy (anonymous):

theta = .5234

OpenStudy (anonymous):

is that \[\pi/2\] ?

OpenStudy (anonymous):

π/2 = 1.57

OpenStudy (anonymous):

im sorry for all the questions but I want to know what the theta you got but in radians